# Dynamical Mean-Field Theory of Complex Systems on Sparse Directed Networks

### v1

@@ Introduction.
Figure 2: Phase diagram $(c,\sigma_{J})$ of the NN model for fixed $10<\mu_{J}<$ 1. The macroscopic variables $m(t)$ and $q(t)$ relax to trivial and nontrivial fixed-points in phases I and II, respectively. In the chaotic phase III, $m(t)$ shows aperiodic oscillations around zero. The dotted line at $c=1$ marks the percolation threshold, while the dashed curve identifies the gap-gapless transition (see the main text). The inset illustrates the sensitivity of $m(t)$ to small perturbations in the initial conditions. Each colored curve represents the numerical solution of Eq. (7) using population dynamics with $5\times 10^{5}$ paths, with initial states sampled from an uniform distribution with unit mean and standard deviation of $\mathcal{O}(10^{-5})$. The red dot is located at $(c^{*},\sigma^{*}_{J})=(\mu_{J}^{-1},\sqrt{\mu_{J}(1-\mu_{J})})$.
We emphasize that the above stability analysis does not give insights into the model’s dynamical behavior. Therefore, we study the dynamics of $m(t)$ and $q(t)$ by solving Eq. (7). Figure 2 presents the resulting phase diagram. In phase I, $m(t)$ and $q(t)$ relax exponentially to the trivial solution $m=q=0$, while in phase II, $m(t)$ and $q(t)$ evolve to nonzero fixed-points. In phase III, the neural network exhibits chaotic behavior, characterized by slow and aperiodic oscillations of $m(t)$ around zero. The inset in Fig. 2 illustrates the sensitivity of $m(t)$ to small perturbations in the initial conditions, a hallmark of deterministic chaos Baker and Gollub (1996). In phase III, the distribution of neural firing rates $\tanh(x_{i}(t))$ does not evolve to a stationary profile. The phase diagram in Fig. 2 is universal, as the transition lines depend only on the first and second moments of $p_{J}$ and $p_{{\rm in},k}$.
Figure 3: Continuous transition between the fixed-point phases for the NN model on directed random networks. The coupling strengths follow a Gaussian distribution with mean $\mu_{J}=1/3$ and standard deviation $\sigma_{J}=0.1$. The results are derived from Eq. (7) using population dynamics with $5\times 10^{5}$ paths. (a) Fixed-point solution $m(t)=m$ as a function of $c$ for Poisson and geometric indegrees (Eq. (9)). (b) Dynamics of $m(t)$ across the transition for Poisson indegrees.
@@
Figure 3 characterizes the transition between phases I and II. As $c$ approaches the critical mean degree $c^{*}=\mu_{J}^{-1}$ from above, the nontrivial fixed-point $m=m(t)$ vanishes continuously. The critical point $c^{*}$ is independent of $p_{{\rm in},k}$, as predicted by Eq. (10), and $m(t)$ relaxes exponentially fast inside phases I and II. Due to critical slowing down, the numerical solution of Eq. (7) becomes computationally more demanding near $c^{*}$. In phase II, the neuronal firing rates evolve to a stationary distribution with a finite variance.
@@
Figure 4: Dynamics of $m(t)$ and $q(t)$ for the NN model on directed random networks with Poisson indegrees. Random coupling strengths follow a Gaussian distribution with mean $\mu_{J}=1/3$ and standard deviation $\sigma_{J}$. The results are derived from the solutions of Eq. (7) using population dynamics with $5\times 10^{5}$ paths. (a) Transition between phases II and III for $c=5$. The critical value of $\sigma_{J}$ is consistent with the dashed line in Fig. 2. (b) Transition between phases I and III for $\sigma_{J}=0.6$. The critical value of $c$ is consistent with Eq. (10).
Figure 4 shows the dynamics of $m(t)$ and $q(t)$ across the transition between the fixed-point phases and the chaotic regime. Since $m(t)$ remains close to zero in both phases I and III, this is not an effective parameter to probe the transition between them. The second moment $q(t)$, instead, clearly identifies the transition between phases I and III, as seen in Fig. 4(b). In this respect, this is analogous to the transition between the paramagnetic and the spin-glass phase Nishimori (2001); Martorell et al. (2024). The critical line between phases II and III is not determined by the above stability analysis. Figure 4(a) illustrates the dynamics of $m(t)$ as the variance $\sigma_{J}^{2}$ of the couplings increases. Notably, the value of $\sigma_{J}$ marking the sudden drop of $m(t)$ and the onset of chaotic oscillations is consistent with the dashed line in Fig. 2 (see Metz for additional results). This strongly suggests that the transition between phases II and III is governed by the gap $|\lambda_{1}-\lambda_{2}|$ in the spectrum of directed random networks.
@@ Conclusions.
We have developed a dynamical mean-field theory of complex systems on sparse directed networks, deriving an exact path-probability equation for the effective dynamics in the limit $N\rightarrow\infty$. Unlike sparse models with symmetric interactions Hatchett et al. (2004); Neri and Bollé (2009), our path-probability equation can be numerically solved using population dynamics Metz et al. (2019). We confirmed the exactness of our general solution by comparing it with numerical simulations of fundamental models in the study of epidemic spreading, neural networks, and ecosystems. Finally, we applied our solution to determine the complete phase diagram of the sparse and directed version of the canonical neural network model of Sompolinsky et al. (1988).
Our work paves the way for exploring the role of sparse heterogeneous networks on the dynamics of ecosystems, coupled oscillators, epidemic spreading, and beyond. Future works include determining the phase diagram of the sparse Lotka-Volterra model Bunin (2017), the influence of the external noise on phase diagrams Schuecker et al. (2018), and the role of network heterogeneities on the critical exponents of complex systems Ferreira and Metz (2023); de Pirey and Bunin (2024). Equation (7) also forms the basis for calculating correlation functions Martí et al. (2018) and deriving approximate dynamical equations for macroscopic order-parameters.
@@ Acknowledgements.
F. L. M. thanks Tuan Minh Pham for many stimulating discussions. The author also thanks Thomas Peron and Tuan Minh Pham for their valuable comments on the manuscript.

### v2

@@ Introduction.
Figure 2: Phase diagrams $(c,\sigma_{J})$ of the NN model for $\mu_{J}=0$, $\mu_{J}\geq 1$, and $10<\mu_{J}<$ 1. The average synaptic current $m(t)$ relaxes to trivial and nontrivial fixed-points in phases I and II, respectively. In the chaotic phases III and IV, $m(t)$ exhibits aperiodic oscillations around zero and nonzero values, respectively (see Fig. 4). The dash-dotted line at $c=1$ marks the percolation threshold, while the dashed curve identifies the gap-gapless transition. Circles and diamonds represent numerical results for the transition lines delimiting phase IV (see the main text), obtained from solutions of Eq. (7) using population dynamics with $5\times 10^{4}$ paths, Poisson indegrees, and two values of $\mu_{J}$ ($\mu_{J}=1/3$ and $\mu_{J}=3/2$). The red dot marks $(c^{*},\sigma^{*}_{J})=(\mu_{J}^{-1},\sqrt{\mu_{J}(1-\mu_{J})})$.
We emphasize that the linear stability analysis of the trivial solution provides no information about the relaxation dynamics or the stationary solutions that emerge when $\boldsymbol{x}=0$ becomes unstable. Therefore, we study the dynamics and the stationary states of the NN model by solving Eq. (7). Figure 2 presents the resulting phase diagrams for different regimes of $\mu_{J}$. In phase I, $m(t)$ relaxes exponentially fast to the trivial solution $m=0$, while in phase II, $m(t)$ evolves to a nonzero fixed-point. In phases III and IV, the neural network exhibits chaotic activity, characterized by slow and aperiodic oscillations of $m(t)$ Sompolinsky et al. (1988); Mastrogiuseppe and Ostojic (2018). The stability lines that delimit phase I are universal, as they depend only on the first and second moments of $p_{J}$ and $p_{{\rm in},k}$.
@@
Figure 3 characterizes the transition between the fixed-point phases I and II. As $c$ approaches the critical mean degree $c^{*}=\mu_{J}^{-1}$ from above, the nontrivial fixed-point $m=m(t)$ vanishes continuously. The critical point $c^{*}$ is independent of $p_{{\rm in},k}$ and $m(t)$ relaxes exponentially fast inside phases I and II. Due to critical slowing down, the numerical solution of Eq. (7) becomes computationally more demanding near $c^{*}$. In phase II, the neuronal firing rates evolve to a stationary distribution with a finite variance.
@@
Figure 3: Continuous transition between the fixed-point phases for the NN model on directed random networks. The coupling strengths follow a Gaussian distribution with mean $\mu_{J}=1/3$ and standard deviation $\sigma_{J}=0.1$. The results are derived from Eq. (7) using population dynamics with $5\times 10^{4}$ paths. (a) Fixed-point solution $m(t)=m$ as a function of $c$ for Poisson and geometric indegrees (Eq. (9)). (b) Relaxation dynamics of $m(t)$ across the transition for Poisson indegrees and initial condition $x_{i}(0)=1$. For $c<\mu_{J}^{-1}$, $m(t)$ relaxes exponentially to $m=0$ as $t\rightarrow\infty$.
Figure 4 shows the dynamics of $m(t)$ within the chaotic phases for several initial conditions. After a transient time $T_{\rm tr}$, $m(t)$ stabilizes into an attractor, oscillating around zero in phase III and around a nonzero value in phase IV. The inset in Fig. 4(a) demonstrates the sensitivity of $m(t)$ to small perturbations in the initial conditions, characteristic of deterministic chaos Baker and Gollub (1996). By solving Eq. (7) and numerically computing the temporal averages Gelbrecht et al. (2020)
[…]
for $T\gg 1$, we estimate the transition lines delimiting phase IV Note (1). For $T\rightarrow\infty$, the transition between phases II and IV is determined by $\Delta$, since $\Delta=0$ in the fixed-point phases and $\Delta>0$ in the chaotic phases. The parameter $M$ distinguishes the chaotic phases: $M=0$ in phase III and $M\neq 0$ in phase IV (see Fig. 4). Our numerical results for the transition between phases III and IV are consistent with the dashed line in Fig. 2, suggesting that this transition is governed by the gap $|\lambda_{1}-\lambda_{2}|$ in the network spectrum. Nevertheless, for $c\rightarrow\infty$ this transition to the chaotic phase III slightly deviates from the gap-gapless transition Mastrogiuseppe and Ostojic (2018), particularly for large $\mu_{J}$.
Figure 4: Dynamics of $m(t)$ inside the chaotic phases for directed networks with Poisson indegrees and a Gaussian distribution $p_{J}$ with mean $\mu_{J}=1/3$. The results follow from the solutions of Eq. (7) using population dynamics with $5\times 10^{4}$ paths and several initial conditions. The dashed lines indicate the value of $M$, Eq. (11), characterizing the chaotic attractor in each phase. (a) $c=2.5$ and $\sigma_{J}=2$. Inset: dynamics of $m(t)$ obtained from the deterministic Eq. (1) for a single network instance ($N=4000$), initial condition $x_{i}(0)=0.15-\epsilon$, and different $\epsilon$. (b) $c=5$ and $\sigma_{J}=0.62$.
@@ Conclusions.
We have developed a dynamical mean-field theory of complex systems on sparse directed networks, deriving an exact path-probability equation for the effective dynamics in the limit $N\rightarrow\infty$. Unlike sparse models with bidirected edges Hatchett et al. (2004); Neri and Bollé (2009), our path-probability equation can be numerically solved using population dynamics Metz et al. (2019). We confirmed the exactness of our general solution by comparing it with numerical simulations of fundamental models in the study of epidemic spreading, neural networks, and ecosystems. Finally, we applied our solution to determine the complete phase diagram of the sparse and directed version of the canonical neural network model of Sompolinsky et al. (1988); Mastrogiuseppe and Ostojic (2018).
The numerical solution of Eq. (7) does not require generating networks from the configuration model Fosdick et al. (2018), providing moderate computational advantages over large-scale simulations of Eq. (1) on networks Silva et al. (2020). Beyond its numerical applications, Eq. (7) provides a foundational framework for studying the stationary states Crisanti and Sompolinsky (2018), computing correlation functions Martí et al. (2018), deriving approximate dynamical equations for macroscopic order-parameters, and developing systematic perturbative approaches Metz et al. (2014); Hertz et al. (2016).
Our work paves the way for exploring the role of sparse heterogeneous networks on the dynamics of ecosystems, coupled oscillators, epidemic spreading, and beyond. Future works include determining the phase diagram of the sparse Lotka-Volterra model Bunin (2017), the influence of external noise on phase diagrams Schuecker et al. (2018), and the role of network heterogeneities on the critical exponents of complex systems Ferreira and Metz (2023); de Pirey and Bunin (2024). Lastly, it would be interesting to connect the present formalism with the cavity method in Tarabolo and Dall’Asta (2024).
@@ Acknowledgements.
F. L. M. thanks Tuan Minh Pham and Alessandro Ingrosso for many stimulating discussions. The author also thanks Thomas Peron and Tuan Minh Pham for their valuable comments on the manuscript. F. L. M. acknowledges support from CNPq (Grant No. 402487/2023-0) and the ICTP through the Associates Programme (2023-2028).